Hi, I am helping my grandson with his homwork and I have no idea how to do algebra.
Equation: x% = 2x + 1 divided by 300
I would appreciate any help you could give me. Thanks

Answer by scott8148(6628)

first, convert the percentage to a decimal number by dividing by 100
__ then solve the equation

x/100=(2x+1)/300 __ multiplying by 300 __ 3x=2x+1 __ subtracting 2x __ x=1

1%=3/300


we grandparents have to stick together...